# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v2

@@ Abstract
We study simultaneous symmetry-breaking in a spontaneous Floquet state, focusing on the specific case of an atomic condensate. We first describe the quantization of the Nambu-Goldstone (NG) modes for a stationary state simultaneously breaking several symmetries of the Hamiltonian by invoking the generalized Gibbs ensemble, which enables a thermodynamical description of the problem. The quantization procedure involves a Berry-Gibbs connection, which depends on the macroscopic conserved charges associated to each broken symmetry and is not invariant under generalized gauge transformations. We extend the formalism to Floquet states simultaneously breaking several symmetries, finding that Goldstone theorem translates into the emergence of Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In the case of a spontaneous Floquet state, there is a genuine temporal FNG mode arising from the continuous time-translation symmetry breaking, whose quantum amplitude provides a rare realization of a time operator in Quantum Mechanics. Furthermore, since they conserve energy, spontaneous Floquet states can be shown to possess a conserved Floquet charge. Both the temporal FNG mode and the Floquet charge are distinctive features of a spontaneous Floquet state, absent in conventional, driven systems. Nevertheless, these also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. We apply our formalism to a particular realization of spontaneous Floquet state, the CES state, which breaks $U(1)$ and time-translation symmetries, representing a time supersolid. Using the Truncated Wigner method, we numerically compute its quantum fluctuations, which are theoretically predicted to be dominated by the temporal FNG mode at long times, observing a remarkable agreement between simulation and theory. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
@@ I Introduction
Noether and Goldstone theorems are two of the most fundamental results in Physics Callen (1985), and can be regarded as the sides of the same coin. The former states that any continuous symmetry is translated into a conservation law, while the latter asserts that solutions spontaneously breaking one of those symmetries have associated a zero-energy mode, known as Nambu-Goldstone (NG) mode. Spontaneous symmetry breaking is still a hot topic of research, motivating the search for novel phases of quantum matter which simultaneously break several symmetries, such as supersolids Léonard et al. (2017); Li et al. (2017), quantum Hall ferromagnets Ezawa (2008); Kharitonov (2012); Kim et al. (2021); Coissard et al. (2022), spinor Stenger et al. (1998); Stamper-Kurn and Ueda (2013); Blinova et al. (2023) and rotating Polkinghorne and Simula (2021) condensates, or quantum droplets Chomaz et al. (2016); C. R. Cabrera and L. Tanzi and J. Sanz and B. Naylor and P. Thomas and P. Cheiney and L. Tarruell (2018); Liu et al. (2019); Zin et al. (2021).
@@
Within the Bogoliubov quasiparticle description, the NG modes emerge as zero-energy modes of the spectrum. Each NG mode is paired with a Gibbs mode, which accounts for the fluctuations of the condensate wavefunction with respect to the associated conserved charge. These fluctuations stem from the fact that a symmetry-broken state cannot be an eigenstate of the charge generating the corresponding symmetry transformation. Remarkably, these Goldstone-Gibbs modes can be elegantly described by a geometrical approach that involves a generalization of the Berry connection defined within a manifold whose variables are the conserved charges, denoted as a result as the Berry-Gibbs connection. This connection is extended to the full manifold which includes the continuous parameters of the broken symmetries, where it becomes a symplectic form that provides the commutation relations between the amplitudes of the Goldstone-Gibbs modes, showing that they behave as conjugate coordinate-momenta. This relation is also translated to the dynamics, where the amplitude of the Gibbs modes behaves as a conserved momentum, since it represents the quantum fluctuations of the corresponding charge, while the amplitude of the Goldstone modes displays a ballistic motion, with a velocity proportional to the momenta. Moreover, by invoking only thermodynamic considerations, we show that the Goldstone-Gibbs sector can be diagonalized in terms of a set of uncoupled NG modes, which determine how the original NG modes hybridize within the different Bogoliubov branches in the low-frequency limit. Our results for simultaneous symmetry breaking in a stationary state thus generalize the original works of Refs. Lewenstein and You (1996); Dziarmaga (2004) to an arbitrary number of broken symmetries, and provide a deeper understanding of the nature of the Goldstone-Gibbs modes.
@@
The generality of the developed formalism allows its application for the study of simultaneous symmetry breaking in Floquet states, where the NG modes are now translated into Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In particular, since spontaneous Floquet states break continuous time-translational symmetry, they present a genuine temporal FNG mode. In combination with the generalized Gibbs ensemble, we develop the $(t,\phi)$ formalism, denoted in this way because of its analogy with the conventional $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012), leading to a thermodynamic description analogous to that of stationary states, which we term as Floquet thermodynamics. For spontaneous Floquet states, this is a direct consequence of energy conservation, which implies the existence of a conserved Floquet charge. Conventional Floquet states also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. This is because driven systems are isoperiodic, i.e., they operate at an externally fixed frequency, which is the conjugate variable of the Floquet charge, in analogy with isobaric and isothermal systems where the environment fixes the pressure or the temperature, respectively.
The quantization of the FNG and Gibbs modes is performed along the same lines as in the stationary case, described by a Berry-Gibbs connection that results in a similar coordinate-momentum correspondence. However, a novel feature is that the amplitude of the temporal FNG mode can be regarded as an effective time coordinate, whose conjugate momentum represents the energy fluctuations. Therefore, the quantum amplitude of the temporal FNG mode provides a unique realization of a time operator in Quantum Mechanics. The temporal FNG mode and the Floquet charge are the distinctive features of a spontaneous Floquet state, absent in driven systems.
@@ II Variational Goldstone Theorem
Time-independent solutions to the equations of motion are given by the equilibrium points $\mathbf{x}_{0}$ which extremize the potential, $\boldsymbol{\nabla}V(\mathbf{x}_{0})=0$. Linearized motion around the equilibrium position is characterized by the eigenvalues $m\Omega^{2}_{i}$ of the matrix $\partial^{2}_{ij}V(\mathbf{x}_{0})$. In particular, potential minima are stable equilibrium points, as then all $\Omega_{i}$ are real and represent the frequencies of the normal modes of oscillation.
@@
A simple example to visualize this behavior is given by a particle in a 2D $y$-independent potential, $V(x,y)=V(x)$, with $V(0)$ a global minimum and $V^{\prime\prime}(0)=m\omega_{x}^{2}$. In that case, if we consider that the particle is placed at the equilibrium position $\mathbf{x}=(0,y_{0})$, small displacements along the $x$-direction lead to oscillations with frequency $\omega_{x}$. However, due to the symmetry of the problem, $\omega_{y}=0$, and any small displacement along the $y$-direction leads to an unbounded motion with constant velocity, $y(t)=y_{0}+v_{y}t$, since there is no restoring force in that direction.
@@
which would yield the exact Schrödinger equation if no restriction were imposed on |$\ket{\Psi(t)}$  [^1] Technically, the extremization condition is $\bra{\delta\Psi(t)}i\hbar\partial_{t}-\hat{H}\ket{\Psi(t)}=0$. The Dirac-Frenkel formalism can also be extended to finite temperatures DeAngelis and Gatoff (1991). Other variational methods are based on a Lagrangian approach, where the trial wave function extremizes some action. A large number of well-known equations fall under this class of approximation, including the Gross-Pitaevskii (GP) equation Pitaevskii and Stringari (2016), the Hartree-Fock (HF) equations Giuliani and Vignale (2005), the Gutzwiller ansatz Jaksch et al. (1998), or the MultiConfiguration Time-Dependent Hartree (MCTDH) method Caillat et al. (2005); Alon et al. (2008). In all cases, if we group the variational parameters in a generic vector $X$, the resulting equations of motion take the form
@@
with $M^{i}_{j}(X)$ some matrix that depends on the background stationary solution. The eigenvalues of $M^{i}_{j}(X)$ are thus the energies of the collective modes. This is indeed the case of the Bogoliubov-de Gennes (BdG) equations for both the GP equation Pitaevskii and Stringari (2016) and the Gutzwiller ansatz Caleffi et al. (2020), and of the Time-Dependent Hartree-Fock Approximation (TDHFA) for the HF equations Thouless (2014).
@@ III Simultaneous symmetry breaking in many-body field theory
In the above equation, $\xi^{\mu},\Delta\Psi$ describe the infinitesimal variations of the coordinates and the field, respectively:
@@
the usual scalar product. The generators $T_{\alpha}$ are Hermitian operators under this scalar product, so symmetry transformations are implemented by unitary operators of the form $U=e^{-i\boldsymbol{\alpha}\cdot\mathbf{T}}$. For instance, for typical symmetries, $T_{\theta}=1$, $T_{\mathbf{x}}=-i\nabla$, $T_{\boldsymbol{\varphi}}=\mathbf{x}\times T_{\mathbf{x}}+\mathbf{S}$.
III.1 Gross-Pitaevskii equation in the generalized Gibbs ensemble
@@
For illustrative purposes, we consider the specific case of a BEC of spin-0 particles close to zero temperature, where the classical scalar field $\Psi$ describes the macroscopic wave function of the condensate. Hereafter, we take the pseudopotential $\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})=g\delta(\mathbf{x}-\mathbf{x}^{\prime})$ as the interacting potential, and set units such that $\hbar=m=gn_{0}=1$, where $n_{0}$ is some characteristic density that rescales the field so it becomes dimensionless, $\Psi\to\sqrt{n_{0}}\Psi$. The classical equation of motion derived from the Lagrangian (10) is then the usual time-dependent GP equation,
with $H_{\rm{GP}}$ the effective nonlinear GP Hamiltonian.
@@
Stationary solutions of the constrained Lagrangian (21) are obtained by extremizing $K$ with respect to variations of the macroscopic wavefunction, leading to the Gross-Pitaevskii-Gibbs (GPG) equation:
@@
We remark that each component $T_{\alpha}\Psi_{0}$ of the vector $\mathbf{T}\Psi_{0}$ is nonvanishing as they are broken symmetries by assumption. Furthermore, we also assume in the following that they are linearly independent, meaning that there is no constant vector $\boldsymbol{\chi}$ satisfying $\chi^{\alpha}T_{\alpha}\Psi_{0}=0$. If this were the case, the conserved charges would not be independent between themselves because then $\chi^{\alpha}Q_{\alpha}=0$. Nevertheless, even in this situation, one could further restrict the vectors $\boldsymbol{\alpha},\mathbf{T},\mathbf{Q}$ in such a way that the components of $\mathbf{T}\Psi_{0}$ are all linearly independent.
The GPG equation (23) still takes the same self-consistent form of Eq. (5). Indeed, the standard time-independent GP equation is retrieved from the GPG equation for $\lambda^{\alpha}=0,~{}\alpha\neq\theta$, and $\lambda^{\theta}=\mu$, where $\mu$ is the chemical potential, the Lagrange multiplier associated to the particle number $N$, emerging due to the spontaneous symmetry-breaking of the $U(1)$-invariance under phase transformations. When additional continuous symmetries are broken, the generalized Gibbs ensemble allows us to treat all of them on equal grounds.
When we insert the stationary GPG wavefunction in the original time-dependent GP equation (20) by taking as initial condition $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x})$, we arrive at
@@
and then take scalar product with $\Psi_{0}$, arriving at
@@
where we note that the conserved charges are macroscopic extensive magnitudes, resulting from integrals over the whole volume. By combining this expression with Eqs. (27), (28), we retrieve the generalization for non-homogeneous systems of the first principle of Thermodynamics (notice that we work at $T=0$),
@@
$\int\mathrm{d}\mathbf{x}~{}\partial_{A}p=Q_{\beta}\partial_{A}\lambda^{\beta}.$ (32)
III.2 BdG equations in the generalized Gibbs ensemble
Once we have derived the spontaneous symmetry-breaking solutions within the generalized Gibbs ensemble, we study the emergence of NG modes in their Bogoliubov spectrum. In a classical context, the BdG equations result from taking $\Psi^{\prime}=\Psi+\varphi$ in the time-dependent GP equation (20) and expanding up to linear order in the field fluctuations $\varphi$,

### v3

@@ Abstract
We study simultaneous symmetry-breaking in a spontaneous Floquet state, focusing on the specific case of an atomic condensate. We first describe the quantization of the Nambu-Goldstone (NG) modes for a stationary state simultaneously breaking several symmetries of the Hamiltonian by invoking the generalized Gibbs ensemble, which enables a thermodynamical description of the problem. The quantization procedure involves a Berry-Gibbs connection, which depends on the macroscopic conserved charges associated to each broken symmetry and whose curvature is not invariant under generalized gauge transformations. We extend the formalism to Floquet states simultaneously breaking several symmetries, where Goldstone theorem translates into the emergence of Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In the case of a spontaneous Floquet state, there is a genuine temporal FNG mode arising from the continuous time-translation symmetry breaking, whose quantum amplitude provides a rare realization of a time operator in Quantum Mechanics. Furthermore, since they conserve energy, spontaneous Floquet states can be shown to possess a conserved Floquet charge. Both the temporal FNG mode and the Floquet charge are distinctive features of a spontaneous Floquet state, absent in conventional, driven systems. Nevertheless, these also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. We apply our formalism to a particular realization of spontaneous Floquet state, the CES state, which breaks $U(1)$ and time-translation symmetries, representing a time supersolid. Using the Truncated Wigner method, we numerically compute its quantum fluctuations, which are theoretically predicted to be dominated by the temporal FNG mode at long times, observing a remarkable agreement between simulation and theory. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
@@ I Introduction
Noether and Goldstone theorems are two of the most fundamental results in Physics, and can be regarded as the sides of the same coin Callen (1985). The former states that any continuous symmetry is translated into a conservation law, while the latter asserts that solutions spontaneously breaking one of those symmetries have associated a zero-energy mode, known as Nambu-Goldstone (NG) mode. Spontaneous symmetry breaking is still a hot topic of research, motivating the search for novel phases of quantum matter which simultaneously break several symmetries, such as supersolids Léonard et al. (2017); Li et al. (2017), quantum Hall ferromagnets Ezawa (2008); Kharitonov (2012); Kim et al. (2021); Coissard et al. (2022), spinor Stenger et al. (1998); Stamper-Kurn and Ueda (2013); Blinova et al. (2023) and rotating Polkinghorne and Simula (2021) condensates, or quantum droplets Chomaz et al. (2016); Cabrera et al. (2018); Liu et al. (2019); Zin et al. (2021).
@@
The NG modes emerge as zero-energy modes of the spectrum of collective excitations. Each NG mode is paired with a Gibbs mode, which accounts for the fluctuations of the quantum state with respect to the associated conserved charge. These fluctuations stem from the fact that a symmetry-broken state cannot be an eigenstate of the charge generating the corresponding symmetry transformation. Remarkably, these Goldstone-Gibbs modes can be elegantly described by a geometrical approach that involves a generalization of the Berry connection defined within a manifold whose variables are the conserved charges, denoted as a result as the Berry-Gibbs connection. This connection is extended to the full manifold which includes the continuous parameters of the broken symmetries, where its curvature gives rise to a symplectic form that provides the commutation relations between the amplitudes of the Goldstone-Gibbs modes, hence behaving as conjugate coordinate-momenta. This relation is also translated to the dynamics, where the amplitude of the Gibbs modes behaves as a conserved momentum, since it represents the quantum fluctuations of the corresponding charge, while the amplitude of the Goldstone modes displays a ballistic motion, with a velocity proportional to the momenta. Moreover, by invoking only thermodynamic considerations, we show that the Goldstone-Gibbs sector can be diagonalized in terms of a set of uncoupled NG modes, which determine how the original NG modes hybridize within the different branches of the spectrum in the low-frequency limit. Our results for simultaneous symmetry breaking in a stationary state thus generalize the original works of Refs. Lewenstein and You (1996); Dziarmaga (2004) to an arbitrary number of broken symmetries, and provide a deeper understanding of the nature of the Goldstone-Gibbs modes.
@@
The generality of the developed framework allows its application for the study of simultaneous symmetry breaking in Floquet states, where the NG modes are now translated into Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In particular, since spontaneous Floquet states break continuous time-translational symmetry, they present a genuine temporal FNG mode. For a complete characterization of the problem, we develop the $(t,\phi)$ formalism, denoted in this way because of its analogy with the conventional $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012). In combination with the generalized Gibbs ensemble, the $(t,\phi)$ formalism leads to a thermodynamic description analogous to that of stationary states, which we term as Floquet thermodynamics. For spontaneous Floquet states, this is a direct consequence of energy conservation, which implies the existence of a conserved Floquet charge. Conventional Floquet states also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. This is because driven systems are isoperiodic, i.e., they operate at an externally fixed frequency, which is the conjugate variable of the Floquet charge, in analogy with isobaric and isothermal systems where the environment fixes the pressure or the temperature, respectively. The temporal FNG mode and the Floquet charge are thus the distinctive features of a spontaneous Floquet state, absent in driven systems.
The quantization of the FNG and Gibbs modes is performed along the same lines as in the stationary case, described by a Berry-Gibbs connection that results in a similar coordinate-momentum correspondence. However, a novel feature is that the amplitude of the temporal FNG mode can be regarded as an effective time coordinate, whose conjugate momentum represents the energy fluctuations. Therefore, the quantum amplitude of the temporal FNG mode represents a unique realization of a time operator in Quantum Mechanics.
@@ II Variational Goldstone Theorem
Time-independent solutions to the equations of motion are given by the equilibrium points $\mathbf{x}_{0}$ which extremize the potential, $\boldsymbol{\nabla}V(\mathbf{x}_{0})=0$. Linearized motion around equilibrium positions is characterized by the eigenvalues $m\Omega^{2}_{i}$ of the matrix $\partial^{2}_{ij}V(\mathbf{x}_{0})$. In particular, potential minima are stable equilibrium points, as then all $\Omega_{i}$ are real and represent the frequencies of the normal modes of oscillation.
@@
A simple example to visualize this behavior is given by a particle in a 2D $y$-independent potential, $V(x,y)=V(x)$, with $V(0)$ a global minimum and $V^{\prime\prime}(0)=m\omega_{x}^{2}$. In that case, if we consider that the particle is placed at the equilibrium position $\mathbf{x}=(0,y_{0})$, small displacements along the $x$-axis lead to oscillations with frequency $\omega_{x}$. However, due to the symmetry of the problem, $\omega_{y}=0$, and any small displacement along the $y$-axis leads to an unbounded motion with constant velocity, $y(t)=y_{0}+v_{y}t$, since there is no restoring force in that direction.
@@
which would yield the exact Schrödinger equation if no restriction were imposed on |$\ket{\Psi(t)}$  [^1] Technically, the extremization condition is $\bra{\delta\Psi(t)}i\hbar\partial_{t}-\hat{H}\ket{\Psi(t)}=0$. The Dirac-Frenkel formalism can also be extended to finite temperatures DeAngelis and Gatoff (1991). Other variational methods are based on a Lagrangian approach, where the trial wave function extremizes some action; indeed, Eq. (3) can be regarded as an effective Lagrangian. A large number of well-known equations fall under this class of approximation, including the Gross-Pitaevskii (GP) equation Pitaevskii and Stringari (2016), the Hartree-Fock (HF) equations Giuliani and Vignale (2005), the Gutzwiller ansatz Jaksch et al. (1998), or the MultiConfiguration Time-Dependent Hartree (MCTDH) method Caillat et al. (2005); Alon et al. (2008). In all cases, if we group the variational parameters in a generic vector $X$, the resulting equations of motion take the form
@@
with $M^{i}_{j}(X)$ some matrix that depends on the background stationary solution. The eigenvalues of $M^{i}_{j}(X)$ are the energies of the collective modes. This is indeed the case of the Bogoliubov-de Gennes (BdG) equations for both the GP equation Pitaevskii and Stringari (2016) and the Gutzwiller ansatz Caleffi et al. (2020), and of the Time-Dependent Hartree-Fock Approximation (TDHFA) for the HF equations Thouless (2014).
@@ III Simultaneous symmetry breaking in many-body field theory
In the above equation, $\xi,\Delta\Psi$ describe the infinitesimal variations of the coordinates and the field under the symmetry transformation, respectively:
@@
the usual scalar product. The generators $T_{\alpha}$ are Hermitian operators under this scalar product, so symmetry transformations are implemented by unitary operators of the form $U=e^{-i\boldsymbol{\alpha}\cdot\mathbf{T}}$. For instance, for typical symmetries, $T_{\theta}=1$, $T_{\mathbf{x}}=-i\nabla$, $T_{\boldsymbol{\varphi}}=\mathbf{x}\times T_{\mathbf{x}}+\mathbf{S}$. We note that these derivations apply for an arbitrary many-body system, including both bosons and fermions.
III.1 Spontaneous symmetry breaking in the generalized Gibbs ensemble
@@
For illustrative purposes, we consider the specific case of a BEC of spin-0 particles close to zero temperature, where the classical scalar field $\Psi$ describes the macroscopic wave function of the condensate. Hereafter, we take the pseudopotential $\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})=g\delta(\mathbf{x}-\mathbf{x}^{\prime})$ as the interacting potential, and set units such that $\hbar=m=gn_{0}=1$, where $n_{0}$ is some characteristic density that rescales the field so it becomes dimensionless, $\Psi\to\sqrt{n_{0}}\Psi$. The condensate dynamics is determined by the classical equation of motion derived from the Lagrangian (10), which is the usual time-dependent GP equation
with $H_{\rm{GP}}$ the effective nonlinear GP Hamiltonian. This equation takes the same self-consistent form of Eq. (4) by choosing $X$ as a two-component vector $X=(\Psi,\Psi^{*})$.
Symmetry-breaking states are described as stationary solutions of the constrained Lagrangian (20), obtained by extremizing $K$ with respect to variations of the macroscopic wavefunction, leading to the Gross-Pitaevskii-Gibbs (GPG) equation:
@@
We remark that each component $T_{\alpha}\Psi_{0}$ of the vector $\mathbf{T}\Psi_{0}$ is nonvanishing as they are broken symmetries by assumption. Furthermore, we also assume in the following that they are linearly independent, meaning that there is no constant vector $\boldsymbol{u}$ satisfying $u^{\alpha}T_{\alpha}\Psi_{0}=0$. If this were the case, the conserved charges would not be independent between themselves because then $u^{\alpha}Q_{\alpha}=0$. Nevertheless, even in this situation, one could further restrict the vectors $\boldsymbol{\alpha},\mathbf{T},\mathbf{Q}$ in such a way that the components of $\mathbf{T}\Psi_{0}$ are all linearly independent.
The standard time-independent GP equation is retrieved from the GPG equation for $\lambda^{\alpha}=0,~{}\alpha\neq\theta$, and $\lambda^{\theta}=\mu$, where $\mu$ is the chemical potential, the Lagrange multiplier associated to the particle number $N$, emerging due to the spontaneous symmetry-breaking of the $U(1)$-invariance under phase transformations. When additional continuous symmetries are broken, the generalized Gibbs ensemble allows us to treat all of them on equal grounds, leading to the GPG equation.
The actual dynamics of the symmetry-breaking state is obtained by inserting the stationary GPG wavefunction in the original time-dependent GP equation (22). Taking as initial condition $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x})$ gives
@@
and then evaluate the scalar product with $\Psi_{0}$, arriving at
@@
where we note that the conserved charges are macroscopic extensive magnitudes, resulting from integrals over the whole volume. By combining this expression with Eqs. (27), (28) and identifying the element of volume as $dV\equiv d\mathbf{x}$, we retrieve the generalization for non-homogeneous systems of the first principle of Thermodynamics (notice that we work at $T=0$),
@@
$\int\mathrm{d}V~{}\partial_{A}p=Q_{\beta}\partial_{A}\lambda^{\beta}.$ (32)
III.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
Once we have characterized the spontaneous symmetry-breaking states within the generalized Gibbs ensemble, we study the emergence of NG modes in their spectrum of excitations. For a condensate, these are described by the BdG equations. In a classical context, the BdG equations result from taking $\Psi^{\prime}=\Psi+\varphi$ in the time-dependent GP equation (22) and expanding up to linear order in the field fluctuations $\varphi$,